Suppose 1 a b 0 р = - 6 0 e f 0 a b 1 с р = 6 0 e f 0 D b 0 с р = 5 1 e f Then 4 D b 5 с р -3 e f

Find a 3 x 3 matrix B where the determinant is 30 and all entries are greater than 0. B=
Answer #1

Solution: Given, 1 a b C 0 e 70 (cf - de) = -6 O a b с = 6 e 1+2 = (-1) (1) (af-be) = 6 (af-be) = -6 (2) 0 b C = 5 R e f 3+12 2 2 Leti B= 13/4 3 4 13/4 2 5 NOW 161 = 2 (15-8) -2 (45-52) +2 (24) = 2 (7) - 2 (42) +262) - 14 - 13 - 1 2 3 = 14 - 26 = 14